Our company's manufacturing overhead (MOH) cost consists of three items: rent, utility, and indirect materials. Rent is a fixed cost. Indirect materials cost is variable (the cost driver is machine hours (MHs)). Utility-is a mixed cost: (i)we pay a fixed monthly fee to our power supplier, and (ii) we make a payment that depends on our usage of utility (the cost driver for the variable component of utility is MHs).

MHs and MOH costs for the last six months of 2018 are below:

MHs Total MOH

July 7,300 $109,900

August 6,200 $101,700

September 10,000 $125,000

October 6,500 $106,400

November 8,500 $128,500

December 5,000 $100,000

December’s total MOH included $44,500 rent and $50,000 utility.

We estimated January 2019 MHs at 7,100. Using the high-low method, our estimated January utility cost is closest to

$ 58,190

$ 64,790

$ 71,000

$ 56,630

$ 110,500
Answer #1

The answer is $58,190 (first option)

Calculations and explanations:


The monthly rent = 44,500 as it is fixed.

Thus overheads excluding monthly rents are:

MOH (excluding rent) MHs
July                              65,400.00      7,300.00
August                              57,200.00      6,200.00
September                              80,500.00    10,000.00
October                              61,900.00      6,500.00
November                              84,000.00      8,500.00
December                              55,500.00      5,000.00

Using the high low method the highest activity level is 10,000 MHs and associated cost = 80500 and lowest is 5,000 MHs and associated cost is 55,500.

Thus variable cost per MH = (80500-55500)/(10000-5000) = $5 per MH. (this is combined variable cost and includes variable utility + indirect material cost).

For December Utility = 50,000. So indirect material cost = 55500-50000 = 5500. Thus indirect material cost per MH = 5500/5000 = $1.1 per MH.

Thus variable rate for utility = 5-1.1 = $3.9 per MH.

So for December 50,000 (total utility) = fixed + (3.9*5000)

or fixed = 50,000 - (3.9*5000)

= $30,500

In January MH = 7100. Thus utility cost = fixed part + variable part

= 30500 + (3.9*7100)

= 30500+27690

= $58,190